Wawa wants to mix cashew nuts worth $6 a pound with peanuts worth $2 a pound to get a 50 pound mixture worth $5 a pound. How much of each is needed
Answer by checkley77(12844)

6x+2(50-x)=5*50
6x+100-2x=250
4x=250-100
4x=150
x=150/4
x=37.5 pouns of $6 cashews.
50-37.5=12.5 pounds of $2 peanuts.
Proof:
6*37.5+2*12.5=250
225+25=250
250=250